A river flows at 5 m/s (average speed) with a volumetric flow rate of 500 m3/s. The river is 30 m above surface of a lake. Determine the maximum power generation potential of the river at that location (units: MW). Ignore frictional losses and turbine efficiency and assume that all of the mechanical energy from the river is used to generate power.

Hint: First use the energy equation to calculate the total mechanical energy of the river at that location per unit mass (in J/kg), and then calculate the maximum power generation potential (1 W = 1 J/s; 1 MW = 106 W).
Answer #1

Mechanical power available = density of water * g * flow rate * height

density of water=1000Kg/m3

Considering no losses

Mechanical power available = 1000*9.81*500*30 = 147.15*106 watt or 147.15 MW Ans